OpenStudy (anonymous):

if \[\log_{\sin x}\cos x-1+\log_{\sin x}\sin 2x \] then the values of x are

OpenStudy (anonymous):

i assume that + means = then cot x = sin2x sinx = +1/root2 , -1/root2

OpenStudy (anonymous):

oh crap! \[\log_{\sin x}\cos x=1+\log_{\sin x}\sin 2x\] this is the Q

OpenStudy (anonymous):

although it doesnt differ much from what you assumed=)

OpenStudy (anonymous):

did u get it now??

OpenStudy (anonymous):

how did you simplify the log terms bro??

OpenStudy (anonymous):

write 1 as log(sinx) to base sinx

OpenStudy (anonymous):

then bring it to the other side..it becomes log(cotx) to base sinx

OpenStudy (anonymous):

ok got it..no need to xplain nemore buddy=)

OpenStudy (anonymous):

thnx